OpenStudy (anonymous):

please simplify step by step (8+i)(2+7i)

OpenStudy (anonymous):

we just cross multiply. to get 16+56i+2i+7i^2. then we just combine like terms to get 16+58i+7i^2

OpenStudy (anonymous):

i thought so too but its wrong..

OpenStudy (anonymous):

the answer is 56i+9 but idk how.

OpenStudy (anonymous):

no that answer is wrong.

OpenStudy (anonymous):

i checked on mathway

OpenStudy (earthcitizen):

\[j ^{2}=-1 \therefore (8+j)(2+j7) = 16+j56+j2-7 = j58-9\]

OpenStudy (anonymous):